### v1

@@ 6.6 Practical Recommendation
The results of our study suggest that the transformer models, and especially MixPL, achieve competitive results with a relatively small number of labeled images, which makes them ideal in a low-data regime. However, the size of the transformer models is significantly larger as compared to the size of the Consistent-Teacher, and thus the resources needed for training these models are also more expensive. Additionally, the models incur higher latency. In very low-data and low-resource regimes, the Consistent-Teacher may be a good alternative.
7 Conclusions and Future Work
A series of experiments were conducted for a systematic evaluation of three state-of-the-art SSOD models (MixPL, Semi-DETR and Consistent-Teacher) on MS-COCO, Pascal VOC and Beetle datasets. Our study explored the models in a practically relevant data regime through a few-shot setting where only a small number of annotated examples are available. Our results reveal that while these models demonstrate some capacity to generalize in low-data scenarios, their performance degrades significantly as the number of labeled instances decreases. Our study is not limited to the exploration of performance but also focuses on practical considerations such as inference time and model size. Overall, our findings provide new insights into the applicability and limitations of current SSOD models in resource-constrained settings.
Future work could explore the behavior of the SSOD and LLM/VLM models for other types of datasets prevalent in various application domains. For example, it is of interest to explore other object detection datasets with a small number of objects and a small number of object instances in an image, or datasets with a small number of objects but a large number of object instances in an image.

### v2

@@ 6.6 Practical Recommendation
The results of our study suggest that the transformer models, and especially MixPL, achieve competitive results with a relatively small number of labeled images, which makes them ideal in a low-data regime. However, the size of the transformer models is significantly larger as compared to the size of the Consistent-Teacher, and thus the resources needed for training these models are also more expensive. Additionally, the models incur higher latency. In very low-data and low-resource regimes, the Consistent-Teacher may be a good alternative. While our study was specifically focused on the comparison of three SSOD frameworks with transformer or CNN-based backbones, we anticipate that the observed trade-offs between accuracy, model size and latency will generalize to other newer SSOD frameworks and architectures, which will likely shift the performance-latency Pareto frontier upwards. However, the relative trend in which transformer-based SSOD methods favor accuracy, while CNN-based methods favor speed is expected to persist.
7 Conclusions, Limitations and Future Work
A series of experiments were conducted for a systematic evaluation of three state-of-the-art SSOD models (MixPL, Semi-DETR and Consistent-Teacher) on MS-COCO, Pascal VOC and Beetle datasets. Our study explored the models in a practically relevant data regime through a few-shot setting where only a small number of annotated examples are available. Our results reveal that while these models demonstrate some capacity to generalize in low-data scenarios, their performance degrades significantly as the number of labeled instances decreases.
Our study is not limited to the exploration of performance but also focuses on practical considerations such as inference time and model size. Overall, our findings provide insights into the applicability and limitations of current SSOD models in resource-constrained settings, by using several datasets that exhibit smaller or larger number of classes and class imbalance. However, scaling our framework to datasets with significantly larger category counts (e.g., LVIS [gupta2019lvis]) or long-tail distributions may present additional challenges.
Future work will explore the behavior of the SSOD models and Large Language Models (LLMs)/Vision Language Models (VLMs) for other types of datasets and distributions prevalent in various application domains. For example, it is of interest to explore other object detection datasets with a small number of objects and a small number of object instances in an image, or datasets with a small number of objects but a large number of object instances in an image, as well as datasets with larger number of objects and long-tail distributions.